ATLAS UK Cloud Support

→ Europe/London
Vidyo

Vidyo

Tim Adye (Science and Technology Facilities Council STFC (GB)), James William Walder (Science and Technology Facilities Council STFC (GB))

Outstanding tickets

  • 148342 UKI-SCOTGRID-GLASGOW less urgent assigned 2020-08-24 11:56:00 UKI-SCOTGRID-GLASGOW with transfer efficiency degraded and many failures

    • One (DPM) disk server with a possible issue, trying to rebuild raidset.
    • Finish current set of drains before tackling above issue
    • JW to send current set of transfer failures to SS.
  • 148339 UKI-NORTHGRID-LIV-HEP less urgent in progress 2020-08-24 08:41:00 0 transfer efficiency to UKI-NORTHGRID-LIV-HEP_LOCALGROUPDISK endpoint

    • No more failures; unclear why endpoint was available given endpoint full.
    • JW to prioritise UK localgroupdisk cleanup.
  • 147979 UKI-NORTHGRID-MAN-HEP less urgent in progress 2020-08-27 05:55:00 UKI-NORTHGRID-MAN-HEP timeout transfer errros and also deletion errors

    • JW updated ticket asking for status report.
  • 146771 UKI-SCOTGRID-ECDF less urgent in progress 2020-08-20 14:44:00 UKI-SCOTGRID-ECDF deletion failures with “The requested service is not available at the moment.”

    • No further changes applied sofar
  • 146651 RAL-LCG2 urgent on hold 2020-08-10 10:59:00 singularity and user NS setup at RAL

    • on hold
  • 146374 UKI-NORTHGRID-SHEF-HEP urgent in progress 2020-07-22 14:53:00 ATLAS pilot jobs idle on UKI-NORTHGRID-SHEF-HEP CE

    • no update
  • 144759 UKI-SCOTGRID-GLASGOW less urgent on hold 2020-08-10 09:54:00 High traffic from UKI-SCOTGRID-GLASGOW on RAL CVMFS Stratum1

    • on hold
  • 142329 UKI-SOUTHGRID-SUSX top priority on hold 2020-06-04 14:05:00 CentOS7 migration UKI-SOUTHGRID-SUSX

    • on hold

CPU

  • RAL

    • Difference in cpu to pledge. updating corepower (set to 10), to 11.7 (approx current value) would give consistent values with internal monitoring.
  • Northgrid

    • General comments on the nature and use of current storage model.
      • Agressive use of storage by the sites; how to move forward, while keeping legacy ‘frameworks’ supported
    • LANCS downtime from disk related issues.
  • London

  • SouthGrid

    • OX back running after Power failure; not yet at full capacity.
  • Scotgrid

Other new issues

  • RAL: Storage downtime; 10hr; batch farm will also be offline.
  • Alessendra email: TPC, DPM, Macaroons
    • ECDF, OX aware of news.

Ongoing issues

  • CentOS7 - Sussex

  • Grand Unified queues

News round-table

  • Dan

    • Migration to new lustre; Atlas scratch disk migration can cause problems (highly volatile).
      • Set up a scratchdisk on another site temporarily ?
      • Could new disk be installed first, and Old data still readable? Probably not.
      • JW to create Jira - ticket, for DDM to advise on the migration.
  • Peter

    • NTR
  • Sam

    • Current storage set at 500TB for CEPH; site to update atlas srr manually to reasonable value.
  • Gareth

    • ntr
  • Tim

    • NTR
  • JW

    • test version of xrootd-ceph plugin with fix to xrdcp -f issue avaialble.

AOB

There are minutes attached to this event. Show them.
    • 10:00 → 10:20
      Status 20m
      • Outstanding tickets 10m

        Outstanding tickets

        • 148342 UKI-SCOTGRID-GLASGOW less urgent assigned 2020-08-24 11:56:00 UKI-SCOTGRID-GLASGOW with transfer efficiency degraded and many failures

          • One (DPM) disk server with a possible issue, trying to rebuild raidset.
          • Finish current set of drains before tackling above issue
          • JW to send current set of transfer failures to SS.
        • 148339 UKI-NORTHGRID-LIV-HEP less urgent in progress 2020-08-24 08:41:00 0 transfer efficiency to UKI-NORTHGRID-LIV-HEP_LOCALGROUPDISK endpoint

          • No more failures; unclear why endpoint was available given endpoint full.
          • JW to prioritise UK localgroupdisk cleanup.
        • 147979 UKI-NORTHGRID-MAN-HEP less urgent in progress 2020-08-27 05:55:00 UKI-NORTHGRID-MAN-HEP timeout transfer errros and also deletion errors

          • JW updated ticket asking for status report.
        • 146771 UKI-SCOTGRID-ECDF less urgent in progress 2020-08-20 14:44:00 UKI-SCOTGRID-ECDF deletion failures with “The requested service is not available at the moment.”

          • No further changes applied sofar
        • 146651 RAL-LCG2 urgent on hold 2020-08-10 10:59:00 singularity and user NS setup at RAL

          • on hold
        • 146374 UKI-NORTHGRID-SHEF-HEP urgent in progress 2020-07-22 14:53:00 ATLAS pilot jobs idle on UKI-NORTHGRID-SHEF-HEP CE

          • no update
        • 144759 UKI-SCOTGRID-GLASGOW less urgent on hold 2020-08-10 09:54:00 High traffic from UKI-SCOTGRID-GLASGOW on RAL CVMFS Stratum1

          • on hold
        • 142329 UKI-SOUTHGRID-SUSX top priority on hold 2020-06-04 14:05:00 CentOS7 migration UKI-SOUTHGRID-SUSX

          • on hold
      • CPU 5m
        • RAL

          • Difference in cpu to pledge. updating corepower (set to 10), to 11.7 (approx current value) would give consistent values with internal monitoring.
        • Northgrid

          • General comments on the nature and use of current storage model.
            • Agressive use of storage by the sites; how to move forward, while keeping legacy ‘frameworks’ supported
          • LANCS downtime from disk related issues.
        • London

        • SouthGrid

          • OX back running after Power failure; not yet at full capacity.
        • Scotgrid

      • Other new issues 5m

        RAL Storage downtime 8th September


        ** htpp-tpc and macaroons
        (Email from Alessandra)
        As you might know from the updates we have given at the TCB in July and August with gsiftp and srm going to be done with http instead. At the moment we are enabling only dcache and DPM sites.

        dcache works well enough with the supported versions with preference with >5.2.18 and 6.2.x previous releases do have minor bugs that require restarting http.
        DPM 1.14 was released yesterday in EPEL this is a required upgrade to enable HTTP-TPC

        Sites will also have to enable macaroons at the moment not all sites have this enabled. Here is an initial list of sites that do not have it, but you can check also Petr's monitoring for the various storage settings. If macaroons is not enabled we will haveto disable TPC for these sites.

        Australia-ATLAS
        FMPhi-UNIBA
        GRIF-LAL
        IFIC-LCG2
        IN2P3-LPC
        IN2P3-LPSC
        NCG-INGRID-PT
        RO-07-NIPNE
        TR-10ULAKBIM
        UKI-SCOTGRID-ECDF
        UKI-SOUTHGRID-OX-HEP
        UNIBE

        Further updates on this ticket ADCINFR-166

        So here are two requirements:

        1) DPM sites need to upgrade to 1.14
        2) dcache sites are ok better is with releases >5.2.18
        3) Make sure macaroons are enabled


        • RAL: Storage downtime; 10hr; batch farm will also be offline.
        • Alessendra email: TPC, DPM, Macaroons
          • ECDF, OX aware of news.

         

    • 10:20 → 10:40
      Ongoing issues 20m
    • 10:40 → 10:50
      News round-table 10m
      • Dan

        • Migration planned to new lustre; Atlas scratch disk migration can cause problems (highly volatile).
          • Set up a scratchdisk on another site temporarily ?
          • Could new disk be installed first, and Old data still readable? Probably not.
          • JW to create Jira - ticket, for DDM to advise on the migration.
      • Peter

        • NTR
      • Sam

        • Current storage set at 500TB for CEPH; site to update atlas srr manually to reasonable value.
      • Gareth

        • ntr
      • Tim

        • NTR
      • JW

        • test version of xrootd-ceph plugin with fix to xrdcp -f issue avaialble.
    • 10:50 → 11:00
      AOB 10m